# Front stability of infinitely steep travelling waves in population biology

### 21 Dec 2023 v1

@@ Abstract
Reaction–diffusion models are often used to describe biological invasion, where populations of individuals that undergo random motility and proliferation lead to moving fronts. Many reaction–diffusion models of biological invasion are extensions of the well–known Fisher–KPP model that describes the spatiotemporal evolution of a 1D population density, $u(x,t)$, as a result of linear diffusion with flux $\mathcal{J}=-\partial u/\partial x$, and logistic growth source term, $S=u(1-u)$. In 2020 Fadai introduced a new reaction–diffusion model of biological invasion with a nonlinear degenerate diffusive flux, $\mathcal{J}=-u\partial u/\partial x$, and the model was formulated as a moving boundary problem on $0<x<L(t)$, with $u(L(t),t)=0$ and $\textrm{d}L(t)/\textrm{d}t=-\kappa u\partial u/\partial x$ at $x=L(t)$ (J Phys A: Math Theor. 53, 095601). Fadai’s model leads to travelling wave solutions with infinitely steep, well–defined fronts at the moving boundary, and the model has the mathematical advantage of being analytically tractable in certain parameter limits. In this work we consider the stability of the travelling wave solutions presented by Fadai for the first time. We aim to provide general insight by first presenting two key extensions of Fadai’s model by considering: (i) generalised nonlinear degenerate diffusion with flux $\mathcal{J}=-u^{m}\partial u/\partial x$ for some constant $m>0$; and, (ii) solutions describing both biological invasion with $\textrm{d}L(t)/\textrm{d}t>0$, and biological recession with $\textrm{d}L(t)/\textrm{d}t<0$. After establishing the existence of travelling wave solutions for these two extensions, our main contribution is to consider stability of the travelling wave solutions by introducing a lateral perturbation of the travelling wavefront. Full 2D time–dependent level–set numerical solutions indicate that invasive travelling waves are stable to small lateral perturbations, whereas receding travelling waves are unstable. These preliminary numerical observations are corroborated through a linear stability analysis that gives more formal insight into short time growth/decay of wavefront perturbation amplitude. Julia–based software, including level-set algorithms, is available on Github to replicate all results in this study.
@@ 1 Introduction
Figure 1: Schematic travelling wave solutions: (a) Fisher–KPP model on $-\infty<x<\infty$; (b) Porous–Fisher model on $-\infty<x<\infty$; (c) Fisher–Stefan model on $x<L(t)$ ; and (d) the Porous–Fisher–Stefan (PFS) model on $x<L(t)$. Arrows on the travelling wave profiles in (a)–(d) indicate that the Fisher–KPP and Porous–Fisher models describe biological invasion only, whereas the Fisher–Stefan and PFS models can describe biological invasion and biological recession. Schematics in (e)–(f) illustrate four possibilities for the linear stability of travelling wave solutions of the PFS model with $u\to 1$ as $x\to-\infty$. The profile at $t=0$ (red) shows the location of the moving boundary where $u=0$ including a transverse perturbation. Profiles at $t_{1}>0$ (green) and $t_{2}>t_{1}>0$ (blue) and shows the evolution of the moving boundary. Profiles in (e)–(f) are stable invading and receding fronts, respectively. Profiles in (g)–(h) are unstable invading and receding fronts, respectively.
In 2010, Du and Lin Du2010 pioneered a different way of overcoming the first limitation of the Fisher–KPP model by re–formulating that model as a moving boundary problem. This approach, sometimes called the Fisher–Stefan model El-Hachem2019 involves working with a linear diffusive flux and logistic source term on $-\infty<x<L(t)$ with $u(L(t),t)=0$ and a Stefan–like moving boundary condition, $\textrm{d}L(t)/\textrm{d}t=-\kappa\partial u/\partial x$ at $x=L(t)$. The constant $\kappa$ is related to the Stefan number in the classical Stefan problem from the heat and mass transfer literature Crank1987 ; Gupta2017 . While moving boundary problems are classically associated with melting and solidification processes in the heat and mass transfer literature BrosaPlanella2019 ; Mitchell2010 , there is an increasing awareness and application of these types of moving boundary problems for biological applications Gaffney1999 ; Shuttleworth2019 ; Dalwadi2020 . In the context of biological invasion, the Fisher–Stefan model has the advantage that it leads to sharp–fronted travelling wave solutions, as illustrated in Figure 1(c), as well las having the flexibility to model both biological invasion with $c>0$, and biological recession with $c<0$ El-Hachem2021 . Therefore, working with a moving boundary model allows us the flexibility to model processes where some initially occupied regions eventually become vacant.
@@
In this work we consider the front stability of travelling wave solutions of the PFS model. Before exploring questions of front stability, we extend Fadai’s travelling wave analysis in two ways by: (i) working with a generalised nonlinear degenerate diffusion model with flux $\mathcal{J}=-u^{m}\partial u/\partial x$ for some constant $m>0$; and (ii) considering solutions describing biological invasion with $\kappa>0$ and $\textrm{d}L(t)/\textrm{d}t>0$, as well as solutions describing biological recession with $\kappa<0$ and $\textrm{d}L(t)/\textrm{d}t<0$. After establishing key properties of travelling wave solutions for this generalised PFS model, we then consider the stability of this broader family of travelling wave solutions by introducing a lateral perturbation of the travelling wavefront, as shown schematically in Figure 1(e)–(h). In summary, 2D time–dependent level–set numerical solutions indicate that models of invasion with $\kappa>0$ are stable to small lateral perturbations, whereas models of biological recession with $\kappa<0$ are unstable. These preliminary numerical observations are corroborated and formalised using linear stability analysis that describes short time growth or decay of the perturbation amplitude. Julia–based software, including the level-set algorithms, is available on Github to replicate all results in this study.
@@ 2.1 Two–dimensional Porous-Fisher-Stefan model
To explore stability of travelling wave solutions of the PFS model we need to re–cast the reaction–diffusion equation on a 2D domain $\Omega(t)$, with a moving boundary, $\partial\Omega(t)$. Within this context the non-dimensional 2D PFS model can be written as
@@
where $U$ is the initial population density, $\mathcal{V}$ is the normal velocity of the interface, $\kappa$ is a constant that determines the speed of the moving boundary, and $m>0$ is the exponent associated with the power law nonlinear diffusivity. Note that Equations (1a)–(1d) are identical to the mathematical model studied by Fadai Fadai2020 when $m=1$ and $u(x,y,t)$ depends only on $x$ and $t$. Therefore, we view Equations (1a)–(1d) as a relatively straightforward extension of Fadai’s model. The main implication of the generalisations introduced here is that the boundary fixing transformation used by Fadai to solve the 1D time–dependent PDE model is not applicable in 2D. Later in Section 2.2 we will describe how to solve this 2D time–dependent PDE model using a level–set method.
Note that the moving boundary condition, Equation (1d), indicates that the front velocity is proportional to $u^{m}\nabla u$, and we have $u=0$ on the moving boundary, $\partial\Omega(t)$. Therefore, if $\mathcal{V}$ is finite, the gradient of $u$ at the moving boundary will be infinite, as illustrated schematically in Figure 1(d) and established mathematically by Fadai Fadai2020 . To deal with this complication numerically we introduce a transformation $\phi=u^{m+1}$ to give
@@ 2.2 Level set method
where $F(x,y,t)$ is the extension velocity field that is defined across $\mathcal{D}$, and satisfies $F=\mathcal{V}$ on $\partial\Omega(t)$. There are several options for choosing $F$. Here we define $F$ by estimating $\mathcal{V}$ at the interface using a standard second-order finite difference approximation and then use orthogonal extrapolation to define $F$ across $\mathcal{D}$ Osher2003 ; Tam2023 .
@@
We solve Equations (4a)–(4e) by discretising $\mathcal{D}$ using a standard uniformly spaced finite difference mesh. Terms on the right-hand side of Equation (4a) are discretised using standard second-order finite-difference stencils. Interpolation is used when $\partial\Omega(t)$ lies between mesh points Tam2022a .
The extension velocity field $F$ is determined by approximating Equation (4e) using a second-order finite difference approximation and orthogonal extrapolation Osher2003 ; Aslam2004 . Given our estimate of $F$ we approximate Equation (4b) using a central scheme Simpson2005 . With these approximation in place we then solve the resulting system of coupled ODEs numerically using the DifferentialEquations package in Julia Rackauckas2017 . This allows us to take advantage of automatic time stepping algorithms within the DifferentialEquations package which means that the key user-defined choice in implementing our level-set method is the size of the mesh used to discretise $\mathcal{D}$.
@@ 3.1 Travelling wave solutions: Numerical explorations
The numerical exploration in Figure 2 extends the kinds of travelling waves considered previously by Fadai who focused on $m=1$ and $\kappa>0$ Fadai2020 . Here we see that different values of $m$ also lead to travelling wave solutions for the PFS model, and that setting $\kappa<0$ leads to receding travelling waves. We now turn to the phase plane to provide mathematical insight into these travelling wave solutions, as well as enabling us to gather additional quantitative information to check the accuracy of our level–set PDE code. It is notable that our estimates of $c$ from the time–dependent PDE model involves specifying $\kappa$ and $m$ as inputs into numerical solution, and we then use the late–time front velocity to provide an estimate of $c$ as an output. This calculation is approximate in the sense that we necessarily introduce truncation error by discretizing $\mathcal{D}$, and then estimate the long–time travelling wave speed $c$ using a finite–time simulation. In Section 3.2 we will use the phase plane provide a more formal argument about the existence of travelling wave solutions, as well as using the associated dynamical system to provide an independent and more accurate estimate of $c$ for each value of $\kappa$ and $m$.
@@ 3.2 Phase plane analysis
which is identical to the system considered by Fadai Fadai2020 when $m=1$. This system is characterised by two equilibrium points: $(0,0)$ and $(1,0)$. It is interesting to note that standard phase plane analysis of travelling wave solutions of the Fisher-KPP model involves establishing the existence of a heteroclinic orbit between these two equilibrium points Murray2002 , however the phase plane analysis is quite different for the PFS model as travelling wave solutions do not correspond to a heteroclinic orbit. Instead, travelling wave solutions of the PFS model correspond to a trajectory that leaves $(1,0)$ and moves through the fourth quadrant of the phase plane to intersect the $\psi$ axis at some point $\psi^{*}<0$ which corresponds to the moving boundary condition. Linearisation confirms that $(1,0)$ is a saddle point for all $c$ and $m$, which means that there is always a trajectory that leaving $(1,0)$ that moves along the unstable manifold into the fourth quadrant of the $(\phi,\psi)$ phase plane.
@@
whose solution, $\psi(\phi)$, directly describes the shape of the trajectories. The relevant solution of Equation (9) must satisfy the associated boundary conditions, $\psi(1)=0$ and $\psi(0)=-c(1+m)/\kappa$. Although we have been unable to find any exact solution for the general case of Equation (9), we are able to make progress analytically for both slowly invading and slowly receding travelling waves. For $|c|\ll 1$, we expand $\psi(\phi)$ as a regular perturbation expansion in $c$, giving $\psi(\phi)=V_{0}(\phi)+cV_{1}(\phi)+\mathcal{O}(c^{2})$. Substituting this expansion into (9) gives:
@@
where $\alpha=36\sqrt{2}-6\sqrt{3}+24\;\mathrm{log}[(\sqrt{3}-1)/(3\sqrt{2}-4)]$.
@@
Figure 4: Relationships between $c$, $\kappa$ and $m$ deriving using the phase plane. Results in (a) show $c$ as a function of $\kappa$ for $m=0.5,1.0,2.0$ in green, blue and red curves, respectively, and the arrow shows the direction of increasing $m$. Results in (a) are obtained by solving Equations (7)–(8) for certain values of $c$ and $m$, and then applying the boundary condition $\kappa=-c(1+m)/\psi^{*}$ to estimate $\kappa$. Results in (b) are for the special case $m=1$ where the solid blue curve is identical to the solid blue curve in (a), and the dashed red curve is the perturbation result, Equation (14). The inset shows a plot that focuses on the dashed rectangular region near $c=0$ where we see that the perturbation result is visually indistinguishable from the numerically–generated phase plane results. Each curve in (a)–(b) obtained using the phase plane is constructed by using 50 equally–spaced values of $c$ for each $m$, and then using the phase plane to solve for $\kappa$. Interpolating the $(c,\kappa)$ values gives the continuous curves in (a)–(b).
For the special case of $m=1$ we can also generate various phase plane trajectories to relate $c$ and $\kappa$ in the same way that we did for the results in Figure 4(a), allowing us to relate $c$ and $\kappa$ for any value of $\kappa$. In addition, for $m=1$ we also have an exact expressions for $c(\kappa)$ (or $\kappa(c)$) relevant for $|c|\ll 1$ given by Equation (14). Results in Figure 4(b) compare our perturbation approximation with the more general relationship from the phase plane where we see that the two curves are visually indistinguishable near $c=0$. Again, it is interesting to note that Equation (14) holds for both slowly invading $c>0$ and slowly receding $c<0$ travelling waves.

### 06 Jan 2024 v2

@@ Abstract
Reaction–diffusion models are often used to describe biological invasion, where populations of individuals that undergo random motility and proliferation lead to moving fronts. Many reaction–diffusion models of biological invasion are extensions of the well–known Fisher–KPP model that describes the spatiotemporal evolution of a 1D population density, $u(x,t)$, as a result of linear diffusion with flux $\mathcal{J}=-\partial u/\partial x$, and logistic growth source term, $S=u(1-u)$. In 2020 Fadai introduced a new reaction–diffusion model of biological invasion with a nonlinear degenerate diffusive flux, $\mathcal{J}=-u\partial u/\partial x$, and the model was formulated as a moving boundary problem on $0<x<L(t)$, with $u(L(t),t)=0$ and $\textrm{d}L(t)/\textrm{d}t=-\kappa u\partial u/\partial x$ at $x=L(t)$ (J Phys A: Math Theor. 53, 095601). Fadai’s model leads to travelling wave solutions with infinitely steep, well–defined fronts at the moving boundary, and the model has the mathematical advantage of being analytically tractable in certain parameter limits. In this work we consider the stability of the travelling wave solutions presented by Fadai for the first time. We aim to provide general insight by first presenting two key extensions of Fadai’s model by considering: (i) generalised nonlinear degenerate diffusion with flux $\mathcal{J}=-u^{m}\partial u/\partial x$ for some constant $m>0$; and, (ii) solutions describing both biological invasion with $\textrm{d}L(t)/\textrm{d}t>0$, and biological recession with $\textrm{d}L(t)/\textrm{d}t<0$. After establishing the existence of travelling wave solutions for these two extensions, our main contribution is to consider stability of the travelling wave solutions by introducing a lateral perturbation of the travelling wavefront. Full 2D time–dependent level–set numerical solutions indicate that invasive travelling waves are stable to small amplitude lateral perturbations, whereas receding travelling waves are unstable. These preliminary numerical observations are corroborated through a linear stability analysis that gives more formal insight into short time growth/decay of wavefront perturbation amplitude. Julia–based software, including level–set algorithms, is available on Github to replicate all results in this study.
@@ 1 Introduction
Figure 1: Schematic travelling wave solutions: (a) Fisher–KPP model on $-\infty<x<\infty$; (b) Porous–Fisher model on $-\infty<x<\infty$; (c) Fisher–Stefan model on $x<L(t)$ ; and (d) the Porous–Fisher–Stefan (PFS) model on $x<L(t)$. Arrows on the travelling wave profiles in (a)–(d) indicate that the Fisher–KPP and Porous–Fisher models describe biological invasion only, whereas the Fisher–Stefan and PFS models can describe biological invasion and biological recession. Schematics in (e)–(f) illustrate four possibilities for the linear stability of travelling wave solutions of the PFS model with $u\to 1$ as $x\to-\infty$. The profile at $t=0$ (red) shows the location of the moving boundary where $u=0$ including a small amplitude transverse perturbation. Profiles at $t_{1}>0$ (green) and $t_{2}>t_{1}>0$ (blue) and shows the evolution of the moving boundary. Arrows in (e)–(h) show the direction of increasing time. Profiles in (e)–(f) are stable invading and receding fronts, respectively. Profiles in (g)–(h) are unstable invading and receding fronts, respectively.
In 2010, Du and Lin Du2010 pioneered a different way of overcoming the first limitation of the Fisher–KPP model by re–formulating that model as a moving boundary problem. This approach, sometimes called the Fisher–Stefan model El-Hachem2019 involves working with a linear diffusive flux and logistic source term on $-\infty<x<L(t)$ with $u(L(t),t)=0$ and a Stefan–like moving boundary condition, $\textrm{d}L(t)/\textrm{d}t=-\kappa\partial u/\partial x$ at $x=L(t)$. The constant $\kappa$ is related to the Stefan number in the classical Stefan problem from the heat and mass transfer literature Crank1987 ; Gupta2017 . While moving boundary problems are classically associated with melting and solidification processes in the heat and mass transfer literature BrosaPlanella2019 ; Mitchell2010 , there is an increasing awareness and application of these types of moving boundary problems for biological applications Gaffney1999 ; Shuttleworth2019 ; Dalwadi2020 ; Jepson2022 . In the context of biological invasion, the Fisher–Stefan model has the advantage that it leads to sharp–fronted travelling wave solutions, as illustrated in Figure 1(c), as well as having the flexibility to model both biological invasion with $c>0$, and biological recession with $c<0$ El-Hachem2021 . Therefore, working with a moving boundary model allows us the flexibility to model processes where some initially occupied regions eventually become vacant.
@@
In this work we consider the front stability of travelling wave solutions of the PFS model. Before exploring questions of front stability, we extend Fadai’s travelling wave analysis in two ways by: (i) working with a generalised nonlinear degenerate diffusion model with flux $\mathcal{J}=-u^{m}\partial u/\partial x$ for some constant $m>0$; and (ii) considering solutions describing biological invasion with $\kappa>0$ and $\textrm{d}L(t)/\textrm{d}t>0$, as well as solutions describing biological recession with $\kappa<0$ and $\textrm{d}L(t)/\textrm{d}t<0$. After establishing key properties of travelling wave solutions for this generalised PFS model, we then consider the stability of this broader family of travelling wave solutions by introducing a small amplitude lateral perturbation of the travelling wavefront, as shown schematically in Figure 1(e)–(h). In summary, 2D time–dependent level–set numerical solutions indicate that models of invasion with $\kappa>0$ are stable to small lateral perturbations, whereas models of biological recession with $\kappa<0$ are unstable. These preliminary numerical observations are corroborated and formalised using linear stability analysis that describes short time growth or decay of the perturbation amplitude. Julia–based software, including the level-set algorithms, is available on Github to replicate all results in this study.
@@ 2.1 Two–dimensional Porous-Fisher-Stefan model
To explore stability of travelling wave solutions of the PFS model we first re–cast the reaction–diffusion equation on a 2D domain $\Omega(t)$, with a moving boundary, $\partial\Omega(t)$. Within this context the non-dimensional 2D PFS model can be written as
@@
where $U$ is the initial population density, $\mathcal{V}$ is the normal velocity of the interface, $\kappa$ is a constant that determines the speed of the moving boundary, and $m>0$ is the exponent associated with the power law nonlinear diffusivity. Note that Equations (1a)–(1d) are identical to the mathematical model studied by Fadai Fadai2020 when $m=1$ and $u(x,y,t)$ depends only on $x$ and $t$. Therefore, we view Equations (1a)–(1d) as a relatively straightforward extension of Fadai’s model. The main implication of the generalisations introduced here is that the boundary fixing transformation used by Fadai to solve the 1D time–dependent PDE model is not applicable in 2D. Later in Section 2.2 we will describe how to solve this 2D time–dependent PDE model using a level–set method Sethian1999 ; Osher2003 ; Tam2023 .
Note that the moving boundary condition, Equation (1d), indicates that the front velocity is proportional to $u^{m}\nabla u$, and we have $u=0$ on the moving boundary, $\partial\Omega(t)$. Therefore, if $\mathcal{V}$ is finite, the gradient of $u$ at the moving boundary will be infinite, as illustrated schematically in Figure 1(d) and established mathematically by Fadai Fadai2020 for $m=1$. To deal with this complication numerically we introduce a transformation $\phi=u^{m+1}$ to give
@@ 2.2 Level set method
where $F(x,y,t)$ is the extension velocity field that is defined across $\mathcal{D}$, and satisfies $F=\mathcal{V}$ on $\partial\Omega(t)$. There are several options for choosing $F$. Here we define $F$ by estimating $\mathcal{V}$ at the interface using a second-order finite difference approximation and then use orthogonal extrapolation to define $F$ across $\mathcal{D}$ Osher2003 ; Tam2023 .
@@
We solve Equations (4a)–(4e) by discretising $\mathcal{D}$ using a standard uniformly spaced finite difference mesh. Terms on the right-hand side of Equation (4a) are discretised using second-order finite-difference approximation using non-constant grid spacing. Interpolation is used when $\partial\Omega(t)$ lies between mesh points Tam2022a .
The extension velocity field $F$ is determined by approximating Equation (4e) using a second-order finite difference approximation and orthogonal extrapolation Osher2003 ; Aslam2004 . Given our estimate of $F$ we approximate Equation (4b) using a central scheme Jiang1998 ; Simpson2005 . With these approximation in place we then solve the resulting system of coupled ODEs numerically using the DifferentialEquations package in Julia Tsitouras2011 ; Rackauckas2017 .
@@ 3.1 Travelling wave solutions: Numerical explorations
The numerical exploration in Figure 2 extends the kinds of travelling waves considered previously by Fadai who focused on $m=1$ and $\kappa>0$ Fadai2020 . Here we see that different values of $m$ also lead to travelling wave solutions for the PFS model, and that setting $\kappa<0$ leads to receding travelling waves. We now turn to the phase plane to provide mathematical insight into these travelling wave solutions, as well as enabling us to gather additional quantitative information to check the accuracy of our level–set PDE code. It is notable that our estimates of $c$ from the time–dependent PDE model involves specifying $\kappa$ and $m$ as inputs into numerical solution, and we then use the late–time front velocity to provide an estimate of $c$ as an output. This calculation is approximate in the sense that we necessarily introduce truncation error by discretizing $\mathcal{D}$, and then estimate the long–time travelling wave speed $c$ using a finite–time simulation. In Section 3.2 we will use the phase plane provide a more formal argument about the existence of travelling wave solutions, as well as using the associated dynamical system to provide an accurate and independent unique value of $c$ for each combination of $\kappa$ and $m$.
@@ 3.2 Phase plane analysis
which is identical to the system considered by Fadai Fadai2020 when $m=1$. This system is characterised by two equilibrium points: $(0,0)$ and $(1,0)$. It is interesting to note that standard phase plane analysis of travelling wave solutions of the Fisher-KPP model involves establishing the existence of a heteroclinic orbit between two equilibrium points Murray2002 , however the phase plane analysis for travelling wave solutions of the PFS model is quite different as travelling wave solutions do not correspond to a heteroclinic orbit in the phase plane. Instead, travelling wave solutions of the PFS model correspond to a trajectory that leaves $(1,0)$ and moves through the fourth quadrant of the phase plane to intersect the $\psi$ axis at some point $\psi^{*}<0$ which corresponds to the moving boundary condition. Linearisation confirms that $(1,0)$ is a saddle point for all $c$ and $m$, and there is always a trajectory leaving $(1,0)$ that moves along the unstable manifold into the fourth quadrant of the $(\phi,\psi)$ phase plane.
@@
whose solution, $\psi(\phi)$, directly describes the shape of the trajectories. The relevant solution of Equation (9) must satisfy $\psi(1)=0$, and then the value of $\kappa$ can be determined by applying $\psi(0)=-c(1+m)/\kappa$. Although we have been unable to find any exact solution for the general case of Equation (9), we are able to make progress analytically for both slowly invading and slowly receding travelling waves. For $|c|\ll 1$, we expand $\psi(\phi)$ as a regular perturbation expansion in $c$, giving $\psi(\phi)=V_{0}(\phi)+cV_{1}(\phi)+\mathcal{O}(c^{2})$. Substituting this expansion into (9) gives:
@@
where $\alpha=36\sqrt{2}-6\sqrt{3}+24\;\mathrm{log}[(\sqrt{3}-1)/(3\sqrt{2}-4)]$ Fadai2020 .
@@
Figure 4: Relationships between $c$, $\kappa$ and $m$ deriving using the phase plane. Results in (a) show $c$ as a function of $\kappa$ for $m=0.5,1.0,2.0$ in green, blue and red curves, respectively, and the arrow shows the direction of increasing $m$. Results in (a) are obtained by solving Equations (7)–(8) for certain values of $c$ and $m$, and then applying the boundary condition $\kappa=-c(1+m)/\psi^{*}$ to estimate $\kappa$. Results in (b) are for the special case $m=1$ where the solid blue curve is identical to the solid blue curve in (a), and the dashed red curve is the perturbation result, Equation (14). The inset shows a plot that focuses on the dashed rectangular region near $c=0$ where we see that the perturbation result is visually indistinguishable from the numerically–generated phase plane results at this scale. Each curve in (a)–(b) obtained using the phase plane is constructed by using 50 equally–spaced values of $c$ for each $m$, and then using the phase plane to solve for $\kappa$. Interpolating the $(c,\kappa)$ values gives the continuous curves in (a)–(b).
For the special case of $m=1$ we can also generate various phase plane trajectories to relate $c$ and $\kappa$ in the same way that we did for the results in Figure 4(a), allowing us to relate $c$ and $\kappa$ for any value of $\kappa$. In addition, for $m=1$ we also have expressions for $c(\kappa)$ (or $\kappa(c)$) relevant for $|c|\ll 1$ given by Equation (14). Results in Figure 4(b) compare our perturbation approximation with the more general relationship from the phase plane where we see that the two curves are visually indistinguishable near $c=0$. Again, it is interesting to note that Equation (14) holds for both slowly invading $c>0$ and slowly receding $c<0$ travelling waves.